# Influence of finite-temperature effects on CMB power spectrum

### v1

@@ Abstract
We explore the implications of finite-temperature quantum field theory effects on cosmological parameters within the framework of the $\Lambda$CDM model and its modification. By incorporating temperature-dependent corrections to the cosmological constant, we extend the standard cosmological model to include additional density parameters, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, which arise from finite-T quantum gravitational effects. Using the Cosmic Linear Anisotropy Solving System, we analyze the impact of these corrections on the cosmic microwave background power spectrum and compare the results with the Planck 2018 data. Through brute-force parameter scans and advanced machine learning techniques, including quartic regression, we demonstrate that the inclusion of $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ improves the model’s predictive accuracy, achieving high $R^{2}$ values and low mean squared error. The present work paves the way for future research into higher-order corrections and enhanced computational methods for cosmological parameter estimation.
@@ 1 Introduction
The conventional hydrodynamic approach as well as the one based on Boltzmann’s equations is essentially a classical framework: it does not take into account the quantum effects introduced by finite-T QFT. In addition to classical thermodynamics, finite-T QFT effects should play a significant role in the cosmological constant, feeding into it through perturbative quantum corrections. Ideally, a second-quantized description whose effective description is provided by the hydrodynamic should be employed. Imagine obtaining the quantum action, i.e., 1PI action. (Only the first several terms in the derivative expansion will matter.) The loop contributions to the CC depend on the temperature. This in turn implies that in the corresponding FRLW-type cosmology the CC comes to be time-dependent through the time-dependence of the temperature [7]. [^1] A time-dependent CC implies the following for the conservation laws [7]. With separate conservation for matter and radiation, one does not impose its separate conservation condition for the CC. Instead, the conservation equation of the total stress tensor follows from the Einstein equation. This bears an interesting implication for renormalization of the Newton’s constant. Finite-T effects were considered earlier [5][6] [^2] These works are finite-temperature extensions of earlier works on foliation-based gravity quantization [8], with one key element being the physical state condition [9] (see [10] for a similar idea). In [11], the gauge $K=0$ (more generally $K=K_{0}$), where $K$ denotes the trace of the second fundamental form, was employed as a crucial ingredient for quantization. A similar approach was later applied in Witten’s recent work [12]. to tackle the long-standing cosmological constant problem. See [18], [19], and [20] for an inspiring review of the CC problem and subsequent works. The CC problem was originally observed from the technical feature of onshell renormalization that the one-loop correction is enormously larger than the observed value of the CC, which necessitated a fine-tuned cancellation between the renormalized value of the CC and the one-loop correction. However, what has been noted in [5, 6, 7] (see also [13] and [14]) was that if one takes the renormalized mass to be on the order of the temperature - which is hinted by the fact that the presence of temperature yields a $T^{4}$-order CC term at the quantum level, the one-loop correction becomes small and this naturally allows one to take a small value of the renormalized value of the CC. It has been proposed that the CC problem may well be a matter of how to manage perturbation theory and finite renormalization, instead of a genuine problem. See also [21] for a recent related discussion.
@@
At the phenomenological level, the crux of our approach is that the finite-T effects naturally introduce additional parameters with which to fit the data. It is not difficult to see how these extra parameter arises. As discussed in the body, the form of the CC after taking the leading finite-T quantum correction into account is $\Lambda=\Lambda_{1}+\Lambda_{2}\frac{1}{a^{4}}+\Lambda_{3}\frac{1}{a^{2}}+\cdots$ where $a$ denotes the scale parameter; $\Lambda_{1},\Lambda_{2}$, and $\Lambda_{3}$ are genuine (viz., time-independent) constants. In view of this it is natural to introduce the corresponding density parameters $\Omega_{\Lambda_{2}},\Omega_{\Lambda_{3}}$ in addition to $\Omega_{\Lambda_{1}}$, the usual vacuum energy density parameter. (One may raise that the parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$ can be absorbed into the $\Omega_{r}$ and $\Omega_{K}$, the radiation and curvature density parameters, respectively; see the discussion below for this.) At the early stage of the matter dominance, the $\Lambda$ term, which is now temperature-dependent and scales as $\sim\frac{1}{a^{4}}$, is not negligible compared with the $a^{-3}$ behavior of matter. In other words, the $\Lambda$ term is a priori expected to make a non-negligible contribution to the Hubble constant between radiation dominance and radiation-matter equality, the period important for some early dark energy models. We thus consider extension of the original $\Lambda$CDM model by including $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$. For the actual analysis of the finite-T modified system, we employ the cosmic linear anisotropy solving system (CLASS) [15, 16].
@@
$(\Omega_{\Lambda_{2}},\;\Omega_{\Lambda_{3}},\;h,\;\omega_{b},\;\omega_{cdm},\;A_{s},\;n_{s},\;\tau_{reio})$ [17], and compare it with the standard $\Lambda$CDM model with $(\Omega_{K},\;h,\;\omega_{b},\;\omega_{cdm},\;A_{s},\;n_{s},\;\tau_{reio})$ in the following manner. For suitable ranges of the parameters, the corresponding power spectra can be generated via CLASS, which can then be plotted to yield the corresponding curve. Each curve being the discrete power spectrum over the multipole moment parameter $\ell$, its Euclidean distance to the Planck experimental curve can therefore be defined. By direct scan it is possible to determine the values of the cosmological parameter set that yields the minimum distance to the Planck 2018 curve. To provide assurance for the result, we also employ statistical and machine learning techniques. While the Planck analysis focused on observational data with advanced Bayesian methods, the present work employs statistical and machine learning regression techniques, specifically quartic regression, to compare the predictive accuracy and generalizability of the two models. Both 7- and 8 parameter models demonstrate exceptional accuracy, with the values of $R^{2}$ (coefficient of determination) nearing 99.9 % for training- as well as testing- dataset. This metric, with mean squared error (MSE), highlights the model’s ability to accurately capture the underlying variance of the distance variable while maintaining low error rates. The results are then compared to the CLASS default model (6 parameters) and the Planck 2018 model (7 parameters).
@@
The rest of the paper is organized as follows. In section 2, we begin by recalling the key results from the work [7] on the finite-temperature quantum gravity (QG) effects. In section 2.1, we adopt Weinberg’s approach [1], essentially reproducing the power spectrum plot derived therein, and compare it with the Planck 2018 results [3]. Although Weinberg’s formalism is not suitable for a detailed comparison with current data, it serves as a useful starting point. This approach aids in understanding the underlying physics at intermediate steps and also provides an analytic expression for the final CMB power spectrum, $\frac{\ell(\ell+1)}{2\pi}C_{\ell}$. Following this, we examine the temperature-corrected cosmological constant (CC). Using Mathematica, we compute the values of the cosmological parameters, modified by the presence of the time-dependent CC, that best fit the fiducial Planck 2018 results. Many of the steps from Chapters 6 and 7 of [1] can be carried over with minimal modifications. For clarity, we demonstrate the analysis by including $\Omega_{2}$ but omitting $\Omega_{3}$ to keep the Mathematica calculations manageable in terms of memory usage. In section 3, we employ a quantum-modified version of the CLASS model [17], which includes the additional parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, and obtain the best fit for the 2018 Planck data [3]. Initially, we perform a brute-force parameter scan, revealing that the theoretical power spectrum is sensitive to both $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, along with the other six parameters. Given the inherent randomness in the sampling method adopted, we validate these results using an independent method. Specifically, we apply various machine learning techniques to assess the predictive performance of both the 7-parameter and 8-parameter models in estimating the distance variable. By employing quartic regression, we optimize the feature values to minimize the predicted distance. The inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model proves to be valuable, as it improves the accuracy of the predictions while maintaining strong generalizability. Finally, in section 4, we conclude with a summary and potential directions for future research.
@@ 2.1 Angular power spectrum in Weinberg’s approach
[…]
@@
[…]
@@
[…]
$\left\{\frac{3(\beta^{2}-1)^{\frac{1}{2}}}{\beta^{4}(1+R_{L})^{\frac{3}{2}}}e^{-2\beta^{2}l^{2}/l_{D}^{2}}{\cal S}^{2}(\beta l/l_{T})\sin^{2}\Big{(}\beta l/l_{H}+\Delta(\beta l/l_{T})\Big{)}\right.$
$\left.+\frac{1}{\beta^{2}(\beta^{2}-1)^{\frac{1}{2}}}\Big{[}3{\cal T}(\beta l/l_{T})R_{L}-(1+R_{L})^{-\frac{1}{4}}e^{-\beta^{2}l^{2}/l_{D}^{2}}{\cal S}(\beta l/l_{T})\cos\Big{(}\beta l/l_{H}+\Delta(\beta l/l_{T})\Big{)}\Big{]}^{2}\right\}.\;$
@@
where the arbitrary pivot scale $k_{\cal R}$ is chose to be $=0.05$, and $d_{T}$, the acoustic horizon distance at last scattering $d_{H}$, and the angular diameter distance $d_{A}$ are given by
@@ 2.2 Finite-temperature modifications
Finite-temperature effects feed the CC through perturbative loops. The CC thus becomes temperature-dependent for a FLRW background. The resulting CC can be viewed as a form of early dark energy that has been present and significant up until around matter-radiation equality. One might mistakenly assume that finite-temperature quantum loop-induced contributions to the cosmological constant, such as those involving photons and neutrinos, are already included as part of the radiation contribution. However, this view overlooks the role of virtual particles in the cosmological constant. The contributions from Standard Model particles (and gravitons) arise from the vacuum energy of virtual particles, while radiation corresponds to the energy density of physical, onshell particles moving through the Universe.
@@
For the analysis in section 3.3, we bring home the distinction between the zero-temperature contributions and the finite-temperature contributions. For the CC problem, the role of the temperature dependence terms was more highlighted [5], [6]. This difference is due to the slight change of the renormalization schemes. Let us recall the renormalization of the cosmological constant proposed in [5] [6], and [7]. It was demonstrated that the one-loop correction is two or three orders of magnitude smaller than the observed value. The essential idea can be explained by considering a gravity-scalar system
$S=\frac{1}{\kappa^{2}}\int d^{4}x\sqrt{-g}\;R-\int d^{4}x\sqrt{-g}\;\Big{(}\frac{1}{2}g^{\mu\nu}\partial_{\mu}\zeta\partial_{\nu}\zeta+V(\zeta)\Big{)}$ (19)
where $\kappa^{2}=16\pi G$ with $G$ being Newton’s constant. The potential $V(\zeta)$ is
$V(\zeta)=\frac{\lambda}{4!}\Big{(}\zeta^{2}+\frac{6}{\lambda}\nu^{2}\Big{)}^{2}.$ (20)
The cosmological constant (CC) value depends, of course, on whether one adopts the complete-square form or the version without the constant term. (Whether to use the complete-square form or the traditional one is not central to the CC problem.) For the onshell potential value, one gets [5] [6]
@@
where the ellipsis includes terms that depend on the mass of the scalar. Let us choose the renormalized mass value appropriately to deal with the CC problem: set the renormalized mass of the scalar particle so that it is on the order of the temperature. At the same time let it be sufficiently smaller than the temperature so that their contributions to the CC can be disregarded. With this renormalization scheme, the leading term in the potential is $\sim T^{4}$, as indicated above. Evaluated at the present temperature of the Universe, this implies that there is a order $10^{4-5}$ gap between the value above and the observed value of the CC, a (potentially) big improvement from the gap noted in the CC problem. As anticipated in [5], once one considers the Standard Model (SM), its particle content should further reduce the gap. To see this, note that a fermionic field contributes
@@
With the role of the leading T-dependent part of the vacuum energy in dealing with the CC problem understood, one can consider the finite-T contributions for the present purpose. (Here we are implicitly employing a new renormalization scheme more convenient for the task at hand; see section 3.3. As we will discuss below in more detail, the motivation for this is to avoid the unnaturalness of the order of $\Omega_{\Lambda_{2}}$, $10^{-8}$.) Let us now consider conducting cosmological perturbation analysis with the 1PI action (with only the first few leading-order terms kept) coupled with the hydrodynamic (or Boltzmann) system. One of the implications of [5, 6, 7] is that the leading quantum corrections yield the following form of the time-dependent vacuum energy [^4] Since higher-order loop corrections are among the terms represented by ()($\cdots$)( ⋯ ), additional time-dependent terms with different $a$-scalings will arise. These terms are not considered in the present work, although they will be addressed in the conclusion.
@@
where $\Lambda_{1},\Lambda_{2}$ are time-independent constants, with $\Lambda_{1}+\Lambda_{2}+\Lambda_{3}+\cdots$ representing the known present value of the CC. The time-dependent CC in eq. (23) introduces additional density parameters, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, corresponding to the $a^{-4}$- and $a^{-2}$-scalings, respectively, in addition to the usual density parameter $\Omega_{\Lambda_{1}}$. Although the CC term associated with $\Omega_{\Lambda_{2}}$ scales as $1/a^{4}$, it does not correspond to radiation, since radiation must satisfy $p=\rho/3$. Similarly, $\Omega_{\Lambda_{3}}$, which appears in the off-shell action through quantum corrections, is conceptually distinct from $\Omega_{K}$, which is associated with the parameter $K$ in the on-shell solution. [^5] At the technical level of the perturbation analysis, however, these two cannot be distinguished. In the analysis presented in section 3, we let $\Omega_{\Lambda_{3}}$ to take on the role of $\Omega_{K}$, setting the latter to zero in the modified CLASS system.
@@
To obtain the CMB power spectrum in terms of the 7 parameters - six usual ones,
$h,\Omega_{B},\Omega_{M},N,n_{S},\tau_{\small{{reio}}},$ (26)
plus $\Omega_{\Lambda_{2}}$, one should first obtain various parameters appearing the angular power spectrum in eq. (8), such as $l_{T},l_{D},l_{H},l_{{\cal R}}$, as functions of the seven parameters. To keep the Mathematica numerical study manageable, for $N,\tau_{\small{{reio}}}$, we adopt the values given in [1], and hone in on the dependence of the various parameters on four cosmological paramters, $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$. Determination of optimal $n_{s}$ can be easily done at the final stage. In other words one can worry about $n_{s}$ at the stage of evaluating $C_{TT,l}^{S}$ after determining $l_{T},l_{D},l_{H},l_{{\cal R}}$ as functions of the independent cosmological parameters: most of our effort is made to obtain the dependences of $l_{T},l_{D},l_{H},l_{{\cal R}}$ on the four parameters $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$. The expressions for $l_{T},l_{D},l_{H},l_{{\cal R}}$ themselves contain parameters, such as $z_{L},\sigma,R_{L},R_{EQ},t_{L}$ (where $\sigma$ denotes the standard deviation of ${\cal O}^{\prime}(T)$ when approximated as a Gaussian distribution), that need to be expressed in terms of the four parameters $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$ in order to quantitatively determine the values of the four parameters for the best fit. Although it would be ideal to obtain these parameters in terms of the full set, $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$, it again turns out that this is a highly memory-demanding procedure and requires more powerful computing resources (this was the case, for instance, for $\sigma$); we will not pursue this full task in the present work. Instead we group $(h,\Omega_{B},\Omega_{M},\Omega_{\Lambda_{2}})$ into $(h,\Omega_{B},\Omega_{\Lambda_{2}})$ and $(h,\Omega_{M},\Omega_{\Lambda_{2}})$, and separately study the dependence of the various parameters on each of these two groups. For each set of $(h,\Omega_{M},\Omega_{\Lambda_{2}})$ and $(h,\Omega_{B},\Omega_{\Lambda_{2}})$, the power spectrum is obtained. Two examples of the plots are given in Fig. 2.666A minor technical note is that the plot of the solid line is that based on CLASS default.ini file but not that of the Planck 2018 data. As we will see in the next section the two are visually indistinguishable. We happend to use the CLASS data since the range of $\ell$ is larger whereas that of Planck 2018 goes up to $\ell=2508$.

### v2

@@ Abstract
[…]
@@ 1 Introduction
The conventional hydrodynamic approach as well as the one based on Boltzmann’s equations is essentially a classical framework: it does not take into account the quantum effects introduced by finite-T QFT. [^1] One might argue that quantum effects are already incorporated through the use of field-theoretic free energy. However, this formulation does not provide an adequate framework for the problem at hand; see Appendix A for a detailed discussion. In addition to classical thermodynamics, finite-T QFT effects should play a significant role in the cosmological constant, feeding into it through perturbative quantum corrections. Ideally, a second-quantized description whose effective description is provided by the hydrodynamics should be employed. Imagine obtaining the quantum action, i.e., 1PI action. (Only the first several terms in the derivative expansion will matter.) The loop contributions to the CC depend on the temperature. This in turn implies that in the corresponding FRLW-type cosmology the CC comes to be time-dependent through the time-dependence of the temperature [7]. [^2] A time-dependent CC implies the following for the conservation laws [7]. (See [8] and the references therein for earlier related works.) With separate conservation for matter and radiation, one does not impose its separate conservation condition for the CC. Instead, the conservation equation of the total stress tensor follows from the Einstein equation. This bears an interesting implication for renormalization of the Newton’s constant. Finite-T effects were considered earlier [5][6] [^3] These works are finite-temperature extensions of earlier works on foliation-based gravity quantization [9], with one key element being the physical state condition [10] (see [11] for a similar idea). In [12], the gauge $K=0$ (more generally $K=K_{0}$), where $K$ denotes the trace of the second fundamental form, was employed as a crucial ingredient for quantization. A similar approach was later applied in Witten’s recent work [13]. to tackle the long-standing cosmological constant problem. See [14], [15], and [16] for an inspiring review of the CC problem and subsequent works. The CC problem was originally observed from the technical feature of onshell renormalization that the one-loop correction is enormously larger than the observed value of the CC, which necessitated a fine-tuned cancellation between the renormalized value of the CC and the one-loop correction. However, what has been noted in [5, 6, 7] (see also [17] and [18]) was that if one takes the renormalized mass to be on the order of the temperature - which is hinted by the fact that the presence of temperature yields a $T^{4}$-order CC term at the quantum level, the one-loop correction becomes small and this naturally allows one to take a small value of the renormalized value of the CC. It has been proposed that the CC problem may well be a matter of how to manage perturbation theory and finite renormalization, instead of a genuine problem. See also [14, 15] and [19] for ealier and more recent related discussions, respectively.
@@
[…]
@@
[…]
It is worth emphasizing that the present approach differs significantly from that of the 2018 Planck collaboration. The latter adheres to the standard Bayesian parameter inference framework: the observed $C_{\ell}$’s are treated as realizations of random variables, while the cosmological parameters are modeled as random variables endowed with prior distributions. This is a well-established Bayesian statistical paradigm. In contrast, the current ML-enhanced deterministic approach treats the cosmological parameters as input variables in a polynomial regression model. (See the introduction of section 3 for the reason and justification for adopting a different approach.) The coefficients of the polynomial - which relate these parameters to the angular power spectrum - are determined through regression. Subsequently, the values of the cosmological parameters that minimize the discrepancy with the 2018 Planck power spectrum curve are identified. Although we are employing frequentist framework, this is not a standard frequentist approach: the cosmological parameters are promoted to input variables in the regression setup, while the coefficients of the polynomial - rather than the cosmological parameters themselves - assume the role of parameters in the statistical sense. The resulting model may be viewed as a surrogate model that approximates the mapping from cosmological parameters to observables. We have more on this in the body.
Both 7- and 8 parameter models demonstrate exceptional accuracy, with the values of the coefficient of determination $R^{2}$ nearing 99.9 % for training- as well as testing- dataset. This metric, with mean squared error (MSE), highlights the model’s ability to accurately capture the underlying variance of the distance variable while maintaining low error rates. The results are then compared to the CLASS default model (6 parameters) and the Planck 2018 model (7 parameters).
@@
The rest of the paper is organized as follows. In section 2, we begin by recalling the key results from the work [7] on the finite-temperature quantum gravity (QG) effects. In section 2.1, we adopt Weinberg’s approach [1], essentially reproducing the power spectrum plot derived therein, and compare it with the Planck 2018 results [3]. Although Weinberg’s formalism is not suitable for a detailed comparison with current data, it serves as a useful starting point. This approach aids in understanding the underlying physics at intermediate steps and also provides an analytic expression for the final CMB power spectrum, $\frac{\ell(\ell+1)}{2\pi}C_{\ell}$. Following this, we examine the temperature-corrected cosmological constant (CC). Using Mathematica, we compute the values of the cosmological parameters, modified by the presence of the time-dependent CC, that best fit the fiducial Planck 2018 results. Many of the steps from Chapters 6 and 7 of [1] can be carried over with minimal modifications. For clarity, we demonstrate the analysis by including $\Omega_{2}$ but omitting $\Omega_{3}$ to keep the Mathematica calculations manageable in terms of memory usage. In Section 3, we begin by highlighting how our approach differs significantly from that of Planck 2018, and we explain the rationale behind this distinction. For computation, we employ a quantum-modified version of the CLASS model [22], which includes the additional parameters $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, and obtain the best fit for the 2018 Planck data [3]. Initially, we perform a brute-force parameter scan, revealing that the theoretical power spectrum is sensitive to both $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, along with the other six parameters. Given the inherent randomness in the subsequent sampling methods adopted, we validate these results using an independent method. Specifically, we apply various machine learning techniques, including random forest and others, to assess the predictive performance of both the 7-parameter and 8-parameter models in estimating the distance variable. We use quartic regression - which performed best among the methods we tested - to optimize the feature values and minimize the predicted distance. The inclusion of $\Omega_{\Lambda_{2}}$ in the 8-parameter model proves to be valuable, as it improves the accuracy of the predictions while maintaining strong generalizability. Finally, in section 4, we conclude with a summary and potential directions for future research. In Appendix A, we distinguish the present framework from quantum field-theoretic thermodynamics formulated through Helmholtz free energy. Specifically, the former is inherently offshell, while the latter is restricted to an onshell description. A summary of the statistical and machine learning terminology used in this work is provided in Appendix B.
@@ 2.1 Angular power spectrum in Weinberg’s approach
${\cal O}=1-\exp\Big[-\int_{t(T)}^{t_{0}}c\sigma_{\cal T}n_{e}dt\Big]=1-\exp\Big[-c\sigma_{T}\int_{T_{0}}^{T}dT^{\prime}\frac{n_{e}(T^{\prime})}{H(T^{\prime})T^{\prime}}\Big]$ (4)
@@
$g(T)=\frac{d}{dT}\exp\Big[-c\sigma_{T}\int_{T_{0}}^{T}dT^{\prime}\frac{n_{e}(T^{\prime})}{H(T^{\prime})T^{\prime}}\Big]=\frac{d}{dT}\exp\Big[-c\sigma_{T}\int_{T_{0}}^{T}dT^{\prime}\frac{n(T^{\prime})X(T^{\prime})}{H(T^{\prime})T^{\prime}}\Big]$ (5)
@@
$\frac{l(l+1)}{2\pi}C_{TT,l}^{S}$ $=$ $\frac{4\pi}{25}T_{0}^{2}N^{2}e^{-2\tau_{\small{{reion}}}}\int_{1}^{\infty}d\beta\left(\frac{\beta l}{l_{{\cal R}}}\right)^{n_{S}-1}$ (8)
$\hskip-65.04256pt\left\{\frac{3(\beta^{2}-1)^{\frac{1}{2}}}{\beta^{4}(1+R_{L})^{\frac{3}{2}}}e^{-2\beta^{2}l^{2}/l_{D}^{2}}{\cal S}^{2}(\beta l/l_{T})\sin^{2}\Big(\beta l/l_{H}+\Delta(\beta l/l_{T})\Big)\right$.
$\left.\hskip-108.405pt+\frac{1}{\beta^{2}(\beta^{2}-1)^{\frac{1}{2}}}\Big[3{\cal T}(\beta l/l_{T})R_{L}-(1+R_{L})^{-\frac{1}{4}}e^{-\beta^{2}l^{2}/l_{D}^{2}}{\cal S}(\beta l/l_{T})\cos\Big(\beta l/l_{H}+\Delta(\beta l/l_{T})\Big)\Big]^{2}\right\}.\;$
@@
where the arbitrary pivot scale $k_{\cal R}$ is chose to be =0.05=0.05, and $d_{T}$, the acoustic horizon distance at last scattering $d_{H}$, and the angular diameter distance $d_{A}$ are given by
@@ 2.2 Finite-temperature modifications
Finite-temperature effects feed the CC through perturbative loops. The CC thus becomes temperature-dependent for a FLRW background. The resulting CC can be viewed as a form of early dark energy that has been present and significant up until around matter-radiation equality. One might mistakenly assume that finite-temperature quantum loop-induced contributions to the cosmological constant, such as those involving photons and neutrinos, are already included as part of the radiation contribution. However, this view overlooks the role of virtual particles in the cosmological constant. The contributions from Standard Model (SM) particles (and gravitons) arise from the vacuum energy of virtual particles, while radiation corresponds to the energy density of physical, onshell particles moving through the Universe. To be specific, let us consider a system described by Boltzmann kinetic theory, such as a photon gas, coupled to a scalar field - e.g., the Higgs field - on a curved spacetime background. The total action is given by
$S=S_{EH}+S_{sf}+S_{kin},$ (19)
where $S_{EH}$ denotes the Einstein-Hilbert action with an appropriately small CC (see the discussion below), $S_{sf}$ is the action for the scalar field, and $S_{kin}$ accounts for the kinetic system. The explicit form of $S_{kin}$ is known in specific cases (see, e.g., [24, 25]), such as for an ideal gas, but will not be required for the present discussion. The corresponding metric field equation contains the well-known expression for the stress-energy tensor:
$T^{\mu\nu}=\frac{1}{\sqrt{-g}}\sum_{n}\delta^{3}(x_{n}(t)-x)\,p_{n}^{\mu}(t)\,p_{n}^{\nu}(t)\,E_{n}(t).$ (20)
To convey the idea without unessential complications, we consider a fixed, non-dynamical background metric. Also, we focus on the $T^{4}$ scaling associated with $\Omega_{\Lambda_{2}}$. The standard treatment focuses solely on the classical thermodynamics of the kinetic system, which yields the familiar exact $T^{4}$ dependence for the energy density, as noted, for example, in footnote 1 of Section 3.1 in [1]. [^4] One might raise that quantum field-theoretic thermodynamics can be conducted by way of introducing free energy. We take up this issue in Appendix A.
In contrast, the present work amounts to examining quantum field-theoretic effects arising from the scalar field sector $S_{sf}$ (and other SM particles and graviton, which do not change the qualitative behavior) at finite temperature. Evidently, such a scalar field cannot be viewed as radiation due to their heavy mass, and they contribute to the CC as virtual particles, i.e., through vacuum bubble diagrams. Those contributions take a form of an infinite series in powers of the temperature, the leading term of which scales as $T^{4}$.
@@
In anticipation of the analysis in section 3.3, let us bring home the distinction between the zero-temperature contributions and the finite-temperature contributions. For the CC problem, the role of the temperature-dependent terms was more highlighted in [5], [6]. This difference is due to the slight change of the renormalization schemes. Let us recall the renormalization of the cosmological constant proposed in [5] [6], and [7]. It was demonstrated that the one-loop correction is two or three orders of magnitude smaller than the observed value. The essential idea can be explained by considering a gravity-scalar system To be specific, let us consider an Einstein-scalar system:
$S_{\text{ren}}$ $=$ $-\int d^{4}x\Bigg[\frac{1}{2}\partial_{\mu}\phi\partial^{\mu}\phi+\frac{1}{2}m^{2}\phi^{2}+\frac{\lambda}{4!}\phi^{4}\Bigg].$ (21)
The cosmological constant (CC) value depends, of course, on whether one adopts the complete-square form or the version without the constant term. (Whether to use the complete-square form or the traditional one is not central to the CC problem.) For the onshell potential value, one gets [29][4] (see Appendix A for more details)
@@
where the ellipsis includes terms that depend on the mass of the scalar. Let us choose the renormalized mass value appropriately to deal with the CC problem: set the renormalized mass of the scalar particle so that it is on the order of the temperature. (See [26] and references therein for earlier works where a CC resolution proposal involves a novel renormalization scheme.) At the same time let it be sufficiently smaller than the temperature so that their contributions to the CC can be disregarded. With this renormalization scheme, the leading term in the potential is $\sim T^{4}$, as indicated above. Evaluated at the present temperature of the Universe, this implies that there is a order $10^{4-5}$ gap between the value above and the observed value of the CC, a (potentially) big improvement from the gap noted in the CC problem. As anticipated in [5], once one considers the Standard Model, its particle content should further reduce the gap. To see this, note that a fermionic field contributes
@@
With the role of the leading T-dependent part of the vacuum energy in dealing with the CC problem understood, one can consider the finite-T contributions for the present purpose. (Here we are implicitly employing a new renormalization scheme more convenient for the task at hand; see section 3.3. As we will discuss there in more detail, the motivation for this is to avoid the unnaturalness of the order of $\Omega_{\Lambda_{2}}$, $10^{-8}$.) Let us now consider conducting cosmological perturbation analysis with the 1PI action (with only the first few leading-order terms kept) coupled with the hydrodynamic (or Boltzmann) system. One of the implications of [5, 6, 7] is that the leading quantum corrections yield the following form of the time-dependent vacuum energy [^6] Since higher-order loop corrections are among the terms represented by $(\cdots)$, additional time-dependent terms with different $a$-scalings will arise. These terms are not considered in the present work, although they will be addressed in the conclusion.
@@
where $\Lambda_{1},\Lambda_{2}$ are time-independent constants, with $\Lambda_{1}+\Lambda_{2}+\Lambda_{3}+\cdots$ representing the known present value of the CC. The time-dependent CC in eq. (24) introduces additional density parameters, $\Omega_{\Lambda_{2}}$ and $\Omega_{\Lambda_{3}}$, corresponding to the $a^{-4}$- and $a^{-2}$-scalings, respectively, in addition to the usual density parameter $\Omega_{\Lambda_{1}}$. Although the CC term associated with $\Omega_{\Lambda_{2}}$ scales as $1/a^{4}$, it does not correspond to radiation, as already discussed. From another angle, note that radiation must satisfy $p=\rho/3$. Similarly, $\Omega_{\Lambda_{3}}$, which appears in the off-shell action through quantum corrections, is conceptually distinct from $\Omega_{K}$, which is associated with the parameter $K$ in the on-shell solution. [^7] At the technical level of the present approximate perturbation analysis, however, these two cannot be distinguished. In the analysis presented in section 3, we let $\Omega_{\Lambda_{3}}$ to take on the role of $\Omega_{K}$, setting the latter to zero in the modified CLASS system.
@@
Weinberg’s approach is insightful and valuable for understanding the physics of key events during the intermediate period between the generation of the CMB and its observation today. However, as noted in his book, its primary aim is not the precise calculation of the power spectrum curve. For such purposes, computational tools like CAMB or CLASS are typically employed.